# Antichiral edge states and Bogoliubov Fermi surfaces in a two-dimensional proximity-induced superconductor

### v2

@@ Abstract
Introduction. The search for Majorana zero modes as emergent excitations in solid-state devices motivated an impressive research activity aimed at realizing the topological superconducting phase. The efforts are mainly oriented to engineer hybrid platforms with ordinary s-wave superconductors in combination with spin-orbit coupling (SOC) and a time-reversal symmetry breaking mechanism, like a magnetic field, magnetic impurities or magnetic materials [1, 2, 3, 4]. Heterostructures of semiconducting InAs and InSb, having a large g-factor and sizable spin-orbit coupling offer a natural playground in this context. Wires fabricated in these materials captured most of the theoretical [5, 6] and experimental early attention. This 1D topological phase has been also explored in engineered two-dimensional configurations [7], and prominent examples are planar Josephson junctions [8, 9, 10, 11, 12]. Beyond the realization of the localized Majorana modes, the study of two-dimensional superconductors with SOC and broken time-reversal symmetry is an active avenue of research [13]: other properties like the superconducting diode effect [14, 15, 16, 17], the formation of Bogoliubov Fermi surfaces [18, 19, 20, 21, 22, 23, 24, 25, 26, 27] and the realization of the Fulde–Ferrell–Larkin–Ovchinnikov state [28] are equally fascinating.
@@
In this work we show the existence of a new topological superconducting phase in a two-dimensional (2D) electron system with SOC when a magnetic field is applied in-plane. The bulk spectrum of this phase is not fully gapped but has a structure of two Dirac-like cones and hosts antichiral Majorana modes that co-propagate along the edges perpendicular to the magnetic field. This feature is in contrast to fully gapped two-dimensional topological superconductors with broken time-reversal symmetry generated by applying a magnetic field perpendicular to the plane, which has Majorana modes propagating clockwise or anticlockwise along the edges [29, 30, 31, 32].
@@
Figure 1: Antichiral edge states: A two-dimensional electron system in proximity with an $s$-wave superconductor is placed in the $x,y$ plane. The Rashba SOC is in the plane as well as the external magnetic field ${\bf B}=(B_{x},B_{y},0)$. A topological phase with antichiral edge states along ${\bf n}_{||}$ exists for $\mu^{2}\leq V^{2}-\Delta_{0}^{2}$ within a range of angles satisfying $|{\bf n}_{||}\cdot{\bf n}_{V}|<\Delta_{0}/V<1$ (see text). (a) and (b) illustrate two different configurations of the magnetic field and the antichiral edge states. The existence of the topological phase is accompanied by the emergence of a single pair of Bogoliubov Fermi Surfaces in the spectrum. (c) BdG spectrum obtained within the lattice model for the magnetic field oriented along ${\bf n}_{x}$. The projected plane-cuts correspond to $(k_{x}=0,k_{y},E)$ and $(k_{x},k_{y}=0,E)$ surfaces.
@@
Topological systems with antichiral states were theoretically predicted in a modified Haldane model [33] and experimental realizations have been reported in photonic systems [34, 35, 36, 37, 38]. Similar co-propagating states were predicted to emerge due to the coupling between electrons and chiral phonons in graphene lattices [39, 40] and models for altermagnets[41]. In contrast to these examples, the edge modes of the two-dimensional superconducting phase we discuss in the present work are akin to lines of Majorana zero modes of stacked topological wires. This type of topological phase, where the boundary modes are present only in specific edges is identified as “weak topological phases” and have been mainly studied in the context of insulators [42] and semimetals [43]. As mentioned in these studies [42], the name does not mean lack of robustness for the edge states. We discuss the spectral properties and the topological characterization of this phase. Very interestingly, it coexists with the development of a single pair of Bogoliubov Fermi surfaces. We focus on parameters similar to those of Al/InAs heterostructures [15, 16] and discuss the experimental signatures of this phase in the current-phase relation (CPR) of Josephson junctions tailored in such 2D platform.
@@
Model. A sketch of the system and configurations we are considering is presented in Fig. 1. We express the Hamiltonian for the 2D electron system with SOC and magnetic field in the basis ${\bf c}_{\bf k}=\left(c_{\bf k,\uparrow},c_{{\bf k},\downarrow}\right)^{T}$. It reads ${\cal H}_{\rm 2D}({\bf k})={\bf c}_{\bf k}^{\dagger}\left[\xi_{\bf k}+H_{\rm SOC}({\bf k})+H_{\rm Z}\right]\;{\bf c}_{\bf k}$, where the first term is the kinetic dispersion relation and the next terms are the SOC and Zeeman Hamiltonians,
@@
A key ingredient for the topological phase we present in this work is a magnetic field with a non-vanishing projection along the direction of the SOC. We hereafter focus on a fully in-plane magnetic field, as indicated in the schemes of Fig. 1(a, b). In what follows we consider ${\bf n}_{V}\equiv{\bf n}_{x}$, which implies breaking the rotational symmetry in the plane. Hence, the spectrum has completely different features along the directions ${\bf n}_{x}$ (parallel) and ${\bf n}_{y}$ (perpendicular) to the magnetic field, respectively. It is important to notice that for $k_{y}=0$, the spin orientation of the SOC is along ${\bf n}_{y}$ and the Hamiltonian Eq. (3) reduces to the model of helical wires originally proposed in Refs. [5, 6]. This model has a topological phase for $\mu^{2}\leq V^{2}-\Delta_{0}^{2}$. Fig. 1(c) illustrates the bulk spectrum of the 2D model within this range of parameters. It can be observed that while the spectrum along $k_{x}$ is gapped, along $k_{y}$ two cones with a quadratic dispersion relation develop close to $k_{y}=0$. They cross zero energy for larger values, leading to the formation of a single pair of Bogoliubov Fermi surfaces. These cones are asymmetric in the sense that one of them intersects the zero-energy plane from below, while the other from above.
@@
Topological properties. The system we consider belongs to the class D of the classification presented in Ref. [44]. In fact, it has time-reversal symmetry broken while preserves particle-hole symmetry. In the case of the 1D system corresponding to $k_{y}=0$, the topological properties are defined by the value of a Berry-Zak phase and by the existence of Majorana zero modes localized at the ends of the wire. In the 2D system we study, the appropriate invariant is similar to that introduced to characterize Dirac semimetals and high-order topology [45, 46, 47, 48, 49, 43, 50, 51]. In order to obtain it, we consider periodic boundary conditions along $x$ and $y$. We calculate the $x$-directed Wilson loop (holonomy) matrix as a function of $k_{y}$, with elements
[…]
where $P$ denotes path ordering, taken over the Brillouin zone of the one-dimensional Hamiltonian with fixed $k_{y}$. The Berry connection is defined as $A^{\ell,\ell^{\prime}}_{k_{x}}(k_{y})=i\langle u^{\ell}_{\bf k}|\partial_{k_{x}}u^{\ell^{\prime}}_{\bf k}\rangle$ and it is calculated from the two lowest-energy eigenstates $\ell$, $\ell^{\prime}$ of the Hamiltonian in Eq. (3). The numerical evaluation of ${\cal W}$ was done by discretizing the path as in Refs. [52, 53]. We define $\theta(k_{y})=\sum_{\ell}\theta_{\ell}(k_{y})$, being $\theta_{\ell}(k_{y})$ the eigenvalues of the Wilson loop [46, 47, 48, 45, 43]. In Fig. 2(b) we show the result for $\theta(k_{y})$ calculated with the parameters which give the spectrum of panel (a). The energies as a function of $k_{y}$ are obtained for a ribbon with $N_{x}=200$ lattice sites with open boundary conditions (OBC) along $x$ and periodic boundary conditions (PBC) along $y$. We can clearly identify the two cones of the bulk states, associated to the formation of a single pair of Bogoliubov Fermi surfaces, confining the range of $k_{y}$ within which $\theta(k_{y})=\pi$. In addition to the bulk states, we identify doubly degenerate edge states depicted in red. They are effectively described by the Hamiltonian
@@
where $\eta^{\dagger}_{\nu,k_{y}}=\eta_{\nu,-k_{y}}$ are Majorana modes that propagate along the $y$ direction with the same velocity $v$, while they are spatially localized at the left and right ($\nu=l,r$) edges of the ribbon. Therefore, these modes are antichiral and include the Majorana zero modes $\eta^{\dagger}_{\nu,0}=\eta_{\nu,0}$ of the topological 1D limit. The velocity is defined by the component of the SOC perpendicular to the magnetic field and we present below an approximate expression on the basis of a low-energy Hamiltonian.
The spatial probability distribution of the lowest-positive-energy antichiral edge states is shown in Fig. 3. In both panels, the ribbon is described with OBC in $x$. Panel (a) shows results obtained with PBC in $y$, where the two edge states confined at the $l$ and $r$ sides of the ribbon can be clearly distinguished. Since, the previous analysis of the topological nature of the edge modes strongly relies on the existence of translational symmetry along the $y$ direction, the question now arises on the robustness of these modes when this symmetry is broken. In panel (b) the spatial probability density is calculated with open boundary conditions in the two directions. The effect of the OBC is to introduce interference effects in the edge states mediated by the bulk states but it does not prevent their formation.
@@
So far, we have considered a fixed orientation of the magnetic field, perpendicular to the ribbon. We would like to stress that the topological phase is not limited to this particular configuration. We recall that the $k_{y}=0$ channel is equivalent to the 1D system and hence, we can rely on the boundaries of the topological phase provided in this limit by Refs. [54, 55, 56, 57, 58] for arbitrary orientations of the magnetic field. The result is
@@
Effective low-energy model. The topological properties, including the nature of the edge states, can be better understood in terms of a low-energy effective Hamiltonian. This can be derived from the original one in Eq. (3) as explained in the Supplementary Material[59]. The result is
@@
being $M({\bf k})=\left(V-E_{\bf k}\right)$, and $\Delta_{x}=\alpha_{R}\Delta_{0}/(2E_{\bf k})$. This representation makes explicit the chiral symmetry ${\cal C}\equiv\tau^{z}$ of the Hamiltonian $H^{{\cal C}}_{\bf k}$, according to which ${\cal C}H^{{\cal C}}({\bf k}){\cal C}^{-1}=-H^{{\cal C}}({\bf k})$. Regarding $k_{y}$ as a parameter, we see that the topological phase of the 1D limit ($k_{y}=0$) extends over a range of values $k_{y}$ satisfying $M(0,k_{y})\geq 0$. We notice that Eq. (S9) is a Jackiw-Rebbi model [60, 61], with an effective mass $M(x,k_{y})$. This model has zero modes at the boundaries if a domain-wall profile is assumed in the $x$-boundary for the mass term. Furthermore, in Ref. [59] we explicitly calculate these edge modes and verify their Majorana nature. The extra term of $H^{{\rm eff}}_{\rm BdG}({\bf k})$ is proportional to $\tau^{0}$ and we can express in the boundary $d^{0}({\bf k})=vk_{y}$ with $v\simeq-\alpha_{\rm R}$. Hence, the eigenstates of the effective Hamiltonian in the boundary are those of $H^{{\cal C}}({\bf k})$ with the energies shifted by $vk_{y}$ as stated in Eq. (5).
@@
Figure 4: Josephson junction: Current-phase relation (CPR) $J(\phi)$ relative to the critical current $J_{c}$ at zero temperature for different magnetic field orientations on the plane with polar angle $\varphi$ (see inset). The two superconductors are modeled by the lattice model with the same parameters as in Fig. 2. The junction has 200 $k_{y}$ channels connected through a row of sites with a tunneling element $t_{\rm J}=t/2$ (see Ref. [59]). The contribution to the CPR arising from the transverse mode with $k_{y}=0$ is shown with dashed lines.
Josephson current. As already discussed in many works, in particular in Refs. [5, 6, 62, 63], the existence of Majorana zero modes in 1D systems generates a peculiar behavior of the current-phase relation (CPR) $J(\phi)$ in Josephson junctions. Namely, because of the $4\pi$-periodicity of the Andreev spectrum, when Majorana zero modes hybridize at the junction, two of these states cross and $J(\phi)$ has a jump at the value of $\phi$ where the crossing takes place. For magnetic fields perpendicular to the SOC, this happens at $\phi=\pi$. For other orientations, the level crossing, hence the jump in the CPR, occurs at different values of $\phi$ [64, 57, 65].
In Josephson junctions of fully gapped 2D topological superconductors with many channels, the hybridization of the propagating Majorana modes can be described by an effective Dirac Hamiltonian with a $\phi$-dependent mass. The Andreev levels associated to the hybridization of these states with a tunneling amplitude ${\cal T}_{\rm J}$ have energies $\varepsilon_{k_{y}}(\phi)=\pm\sqrt{(vk_{y})^{2}+{\cal T}_{\rm J}^{2}\cos^{2}(\phi/2)}$ [66]. The corresponding contribution to the CPR has a jump at $k_{y}=0$ as in the case of 1D systems, while the other edge-channels have a smooth non-sinusoidal behavior. In contrast, the other $k_{y}$ channels (those not corresponding to the propagating edge states, i.e. the bulk modes) have a dispersion relation of the form $\varepsilon_{k_{y}}(\phi)=\pm E_{{\rm J},k_{y}}\cos(\phi)$, being $E_{{\rm J},k_{y}}$ a characteristic Josephson energy. This is similar to non-topological superconductors and contributes to the CPR as $\propto\sin(\phi)$. The total $J(\phi)$, results from the contribution of all the Andreev states with negative energy. Therefore, the overall shape and features of the CPR are determined by the relative spectral weight of the edge modes among the transverse channels, and by the degree of their hybridization compared to the transparency for the bulk states.
@@
The behavior of the CPR in the topological superconductor with antichiral Majorana modes is illustrated in Fig. 4. The configuration is shown in the inset and details on the calculations are presented in Ref. [59]. The function $J(\phi)$ relative to the critical current for 200 $k_{y}$-channels is shown for different angles $\varphi$ between ${\bf n}_{x}$ and ${\bf n}_{V}$. The supercurrent contribution associated to the zero mode $k_{y}=0$ is shown in dashed lines, where we can clearly identify the jump at $\phi=\pi$ when $\varphi=0$. We see that for $\varphi\neq 0$, while satisfying the condition of Eq. (6), the discontinuity in the supercurrent contribution corresponding to $k_{y}=0$ is shifted to $\phi\neq\pi$. Taking into account the other $k_{y}$-channels, the impact of the zero mode is also visible within the 2D topological phase. For orientations $\varphi$ beyond the condition of Eq. (6) for the boundary of the topological phase, the response is sinusoidal for all the modes as in a non-topological superconductor (grey curve). In all the cases, a response with $J(\phi=0)\neq 0$ is observed for $\varphi\neq 0$, typical of the anomalous Josephson effect. The amplitude of the critical current is also different for different signs of $\phi$, which is the characteristic feature of the superconducting diode effect [15, 16]. Similar behavior has been discussed in the framework of 1D topological superconductors [57, 65].
@@
Conclusions. We have shown the existence of a new topological phase in a 2D proximity-induced superconductor with Rashba spin-orbit coupling. This phase develops when a magnetic field is applied in-plane and is characterized by antichiral Majorana modes propagating along the edges perpendicular to the magnetic field. This phase takes place for a range of parameters (chemical potential and magnetic field) where the Bogoliubov-de Gennes spectrum has a single pair of Bogoliubov Fermi surfaces. We have shown that signatures of these modes can be identified in the behavior of the current-phase relation of a wide Josephson junction with the edges states oriented as sketched in Fig. 4. The topological phase persists for a range of angles close to this configuration.
@@
The experimental realization of this phase is possible in various material platforms. It could be explored in hybrid superconductor/semiconductor heterostructures if a strategy is developed to control the chemical potential of the 2DEG. This could be realized using two quantum wells, one of them buried deep in the heterostructure to be used as a back-gate and the other one very shallow to be proximitized by the thin superconducting film. For Al/InAs heterostructures, for example, high-transparency Josephson junctions have been demonstrated, exhibiting anomalous Josephson effect and diode effect, while supporting large in-plane magnetic fields ($\sim$1.5T) [15, 16]. The gate tunability would allow to study the progressive effect of the emergence of pairs of Bogoliubov Fermi Surfaces at reasonable fields until the conditions for the topological phase are reached. Another potential platform is a two-dimensional magnetic topological insulator like that studied in Ref. [28].
@@
I Details of the derivation of the effective low-energy Hamiltonian

### v3

@@ Abstract
Introduction. The search for Majorana zero modes in solid-state devices has driven intense efforts to realize topological superconductivity. Most approaches engineer hybrid platforms combining conventional s-wave superconductors, spin-orbit coupling (SOC) and time-reversal-symmetry breaking mechanism, like a magnetic field, magnetic properties of materials [Alicea2012Jun, Prada2020Oct, Flensberg2021Oct, Schiela2024Sep]. Heterostructures of semiconducting InAs and InSb, having a large g-factor and sizable spin-orbit coupling offer a natural playground in this context. Wires fabricated in these materials captured most of the theoretical [Lutchyn2010Aug, Oreg2010Oct] and experimental early attention. This topological phase has been also explored in engineered two-dimensional (2D) configurations [Kjaergaard2016Sep], and prominent examples are planar Josephson junctions [Pientka2017May, Hart2017Jan, Fornieri2019May, Ren2019May, Banerjee2023Jun]. Beyond the realization of the localized Majorana modes, 2D superconductors with SOC and broken time-reversal symmetry also exhibit rich physics [Moehle2021Dec], like the superconducting diode effect [Ando2020Aug, BaumgartnerPRL2021, BaumgartnerNature2022, Banerjee2023Nov], the formation of Bogoliubov Fermi surfaces [Agterberg2017Mar, Yuan2018Mar, setty2020bogoliubov, Shaffer2020Jun, Dutta2021Sep, Phan2022Mar, Babkin2024May, Mateos2024Aug, Pal2024Dec, Cohen2024Apr, Ohashi2024Sep] and the realization of the Fulde–Ferrell–Larkin–Ovchinnikov state [Mandal2024Jun] are equally fascinating.
@@
In this work we show the existence of a new topological superconducting phase in a 2D electron system with SOC when a magnetic field is applied in-plane. The bulk spectrum of this phase is not fully gapped but has a structure of two Dirac-like cones and hosts antichiral Majorana modes that co-propagate along the edges perpendicular to the magnetic field. This feature is in contrast to fully gapped 2D topological superconductors with broken time-reversal symmetry generated by applying a magnetic field perpendicular to the plane, which has Majorana modes propagating clockwise or anticlockwise along the edges [Sau2010Jan, Alicea2010Mar, Bernevig2013Apr, Sato2017May]. Such phase is characterized by a Chern number and has been found to be robust against a tilt of the magnetic field with a finite in-plane projection [Loder2015Oct].
@@
Figure 1: Antichiral edge states: A 2D electron system in proximity with an $s$-wave superconductor is placed in the $x,y$ plane. The Rashba SOC is in the plane as well as the external magnetic field ${\bf B}=(B_{x},B_{y},0)$. A topological phase with antichiral edge states along ${\bf n}_{||}$ exists for $\mu^{2}\leq V^{2}-\Delta_{0}^{2}$ within a range of angles satisfying $|{\bf n}_{||}\cdot{\bf n}_{V}|<\Delta_{0}/V<1$ (see text). (a) and (b) illustrate two different configurations of the magnetic field and the antichiral edge states. The existence of the topological phase is accompanied by the emergence of a single pair of Bogoliubov Fermi Surfaces in the spectrum. (c) BdG spectrum obtained within the lattice model for the magnetic field oriented along ${\bf n}_{x}$. The projected plane-cuts correspond to $(k_{x}=0,k_{y},E)$ and $(k_{x},k_{y}=0,E)$ surfaces.
@@
Topological systems with antichiral states were theoretically predicted in a modified Haldane model [Colomes2018Feb] and experimental realizations have been reported in photonic systems [Zhou2020Dec, Xi2023Apr, Wang2021Feb, Chen2022Jun, Hu2024Nov]. Similar co-propagating states were predicted to emerge due to the coupling between electrons and chiral phonons in graphene lattices [MedinaDuenas2022Feb, Mella2023Dec] and models for altermagnets[Sorn2025Apr]. In contrast to these examples, the edge modes of the 2D superconducting phase we discuss in the present work are akin to lines of Majorana zero modes of stacked topological wires. This type of topological phase, where the boundary modes are present only in specific edges is identified as “weak topological phases” and have been mainly studied in the context of insulators [Ringel2012Jul] and semimetals [Wieder2020Jan]. As mentioned in these studies [Ringel2012Jul], the name does not mean lack of robustness for the edge states. We discuss the spectral properties and the topological characterization of this phase. Very interestingly, it coexists with the development of a single pair of Bogoliubov Fermi surfaces. We focus on parameters similar to those of Al/InAs heterostructures [BaumgartnerPRL2021, BaumgartnerNature2022] and discuss the experimental signatures of this phase in the current-phase relation (CPR) of Josephson junctions tailored in such 2D platform.
@@
Model. A sketch of the system is presented in Fig. 1. We express the Hamiltonian for the 2D electron system with SOC and magnetic field in the basis ${\bf c}_{\bf k}=\left(c_{\bf k,\uparrow},c_{{\bf k},\downarrow}\right)^{T}$. It reads ${\cal H}_{\rm 2D}({\bf k})={\bf c}_{\bf k}^{\dagger}\left[\xi_{\bf k}+H_{\rm SOC}({\bf k})+H_{\rm Z}\right]\;{\bf c}_{\bf k}$, where the first term is the kinetic dispersion relation and the next terms are the SOC and Zeeman Hamiltonians,
@@
A crucial ingredient for the topological phase we study is a magnetic field with a non-vanishing projection along the direction of the SOC. We hereafter focus on a fully in-plane magnetic field, as indicated in the schemes of Fig. 1(a, b). In what follows we consider ${\bf n}_{V}\equiv{\bf n}_{x}$, which implies breaking the rotational symmetry in the plane. Hence, the spectrum has completely different features along the directions ${\bf n}_{x}$ (parallel) and ${\bf n}_{y}$ (perpendicular) to the magnetic field, respectively. It is important to notice that for $k_{y}=0$, the spin orientation of the SOC is along ${\bf n}_{y}$ and the Hamiltonian Eq. (3) reduces to the model of helical wires originally proposed in Refs. [Lutchyn2010Aug, Oreg2010Oct]. This model has a topological phase for $\mu^{2}\leq V^{2}-\Delta_{0}^{2}$. Fig. 1(c) illustrates the bulk spectrum of the 2D model within this range of parameters. It can be observed that while the spectrum along $k_{x}$ is gapped, along $k_{y}$ two cones with a quadratic dispersion relation develop close to $k_{y}=0$. They cross zero energy for larger values, leading to the formation of a single pair of Bogoliubov Fermi surfaces, as already pointed out in Refs. [Yuan2018Mar, Babkin2024May]. These cones are asymmetric in the sense that one of them intersects the zero-energy plane from below, while the other from above.
@@
Topological properties. The system we consider belongs to the class D of the classification presented in Ref. [Schnyder2009May]. In fact, it has time-reversal symmetry broken while preserves particle-hole symmetry. In the case of the 1D system corresponding to $k_{y}=0$, the topological properties are defined by the value of a Berry-Zak phase and by the existence of Majorana zero modes localized at the ends of the wire. In the 2D system we study, the appropriate invariant is similar to that introduced to characterize Dirac semimetals and high-order topology [Fidkowski2011Jul, Alexandradinata2014Apr, Cano2018Jun, Bradlyn2019Jan, Bouhon2019Nov, Wieder2020Jan, Wang2019Sep, Awoga2022Apr]. We consider periodic boundary conditions along $x$ and $y$ and calculate the $x$-directed Wilson loop (holonomy) matrix with elements
${\cal W}^{\ell,\ell^{\prime}}(k_{y})=P\exp\left\{i\int_{0}^{2\pi}dk_{x}A^{\ell,\ell^{\prime}}_{k_{x}}(k_{y})\right\},$ (4)
where $P$ denotes path ordering taken over the Brillouin zone of the one-dimensional Hamiltonian with fixed $k_{y}$. This generalizes the Zak phase along $x$ for a fixed value of $k_{y}$. In the present case, the Berry connection is non-abelian and is defined as $A^{\ell,\ell^{\prime}}_{k_{x}}(k_{y})=i\langle u^{\ell}_{\bf k}|\partial_{k_{x}}u^{\ell^{\prime}}_{\bf k}\rangle$. The indices $\ell$, $\ell^{\prime}$ label the two lowest-energy eigenstates of the Hamiltonian in Eq. (3). We evaluate $\cal W$} by discretizing the path as in Refs. [Ortiz1996Nov, PerezDaroca2021Sep], then diagonalize it to obtain the eigenvalues $\theta_{\ell}(k_{y})$. The phase $\theta(k_{y})=\sum_{\ell}\theta_{\ell}(k_{y})=\pi,0$ defines the topological characterization[Alexandradinata2014Apr, Cano2018Jun, Bradlyn2019Jan, Fidkowski2011Jul, Wieder2020Jan]. This, respectively, indicates the existence or non existence of edge states for that $k_{y}$. In Fig. 2(b) we show the result for $\theta(k_{y})$ calculated with the parameters which give the spectrum of panel (a). The energies as a function of $k_{y}$ are obtained for a ribbon with $N_{x}=200$ lattice sites with open boundary conditions (OBC) along $x$ and periodic boundary conditions (PBC) along $y$. We can clearly identify the two cones of the bulk states, associated to the formation of a single pair of Bogoliubov Fermi surfaces, confining the range of $k_{y}$ within which $\theta(k_{y})=\pi$. In addition to the bulk states, we identify doubly degenerate edge states depicted in red. They are effectively described by the Hamiltonian
@@
where $\eta^{\dagger}_{\nu,k_{y}}=\eta_{\nu,-k_{y}}$ are Majorana modes that propagate along the $y$ direction with the same velocity $v$, while they are spatially localized at the left and right ($\nu=l,r$) edges of the ribbon. Therefore, these modes are antichiral and include the Majorana zero modes $\eta^{\dagger}_{\nu,0}=\eta_{\nu,0}$ of the topological 1D limit. The velocity is defined by the component of the SOC perpendicular to the magnetic field and we present below an approximate expression on the basis of a low-energy Hamiltonian. Akin to the models of photonic systems [Colomes2018Feb, Zhou2020Dec, Chen2022Jun], these modes transport energy in one way along the edges, while there is a counterflow through the bulk states.
The spatial probability distribution of the lowest-positive-energy antichiral edge states is shown in Fig. 3. In both panels, the ribbon is described with OBC in $x$. Panel (a) shows results obtained with PBC in $y$, where the two edge states confined at the $l$ and $r$ sides of the ribbon can be clearly distinguished. Because this analysis assumes translational invariance along the $y$ direction, the question now arises on the robustness of these modes when this symmetry is broken. In panel (b) the spatial probability density is calculated with open boundary conditions in the two directions. The effect of the OBC is to introduce interference effects in the edge states mediated by the bulk states but it does not prevent their formation. Similar features are observed in a system with circular shape (see Ref. [sm]).
@@
So far, we have considered a fixed orientation of the magnetic field, perpendicular to the ribbon. We would like to stress that the topological phase is not limited to this particular configuration. We recall that the $k_{y}=0$ channel is equivalent to the 1D system and hence, we can rely on the boundaries of the topological phase provided in this limit by Refs. [Osca2014Jun, Rex2014Sep, Klinovaja2015Mar, Aligia2020Dec, Gruneiro2023Jul] for arbitrary orientations of the magnetic field. The result is
@@
Effective low-energy model. The topological properties, including the nature of the edge states, can be better understood in terms of the following low-energy effective Hamiltonian (see Ref. [sm]),
@@
being $M({\bf k})=\left(V-E_{\bf k}\right)$, and $\Delta_{x}=\alpha_{R}\Delta_{0}/(2E_{\bf k})$. This representation makes explicit the chiral symmetry ${\cal C}\equiv\tau^{z}$ of the Hamiltonian $H^{{\cal C}}_{\bf k}$, according to which ${\cal C}H^{{\cal C}}({\bf k}){\cal C}^{-1}=-H^{{\cal C}}({\bf k})$. Regarding $k_{y}$ as a parameter, we see that the topological phase of the 1D limit ($k_{y}=0$) extends over a range of values $k_{y}$ satisfying $M(0,k_{y})\geq 0$. We notice that Eq. (S9) is a Jackiw-Rebbi model [Jackiw1976Jun, Goldstone1981Oct], with an effective mass $M(x,k_{y})$. This model has zero modes at the boundaries if a domain-wall profile is assumed in the $x$-boundary for the mass term. Furthermore, in Ref. [sm] we explicitly calculate these edge modes and verify their Majorana nature. The extra term of $H^{{\rm eff}}_{\rm BdG}({\bf k})$ is proportional to $\tau^{0}$ and we can express in the boundary $d^{0}({\bf k})=vk_{y}$ with $v\simeq-\alpha_{\rm R}$. Hence, the eigenstates of the effective Hamiltonian in the boundary are those of $H^{{\cal C}}({\bf k})$ with the energies shifted by $vk_{y}$ as stated in Eq. (5).
@@
Figure 4: Josephson junction: Current-phase relation (CPR) $J(\phi)$ relative to the critical current $J_{c}$ at zero temperature for different magnetic field orientations on the plane with polar angle $\varphi$ (see inset). The two superconductors are modeled by the lattice model with the same parameters as in Fig. 2. The junction has 200 $k_{y}$ channels connected through a row of sites with a hopping $t_{\rm J}=t/2$ (see Ref. [sm]). The contribution to the CPR arising from the transverse mode with $k_{y}=0$ is shown with dashed lines.
@@
Josephson current. The behavior of the the current-phase relation (CPR) $J(\phi)$ in the topological superconductor with antichiral Majorana modes is illustrated in Fig. 4. The inset shows the configurations for different angles $\varphi$ between the junction (${\bf n}_{||}$) and the magnetic field (${\bf n}_{V}$). Calculations were done considering 200 $k_{y}$-channels and details on the calculations are presented in Ref. [sm]. For $\varphi$ within the topological phase (satisfying the condition of Eq. (6)), the supercurrent contribution associated to the zero mode $k_{y}=0$ is shown in dashed lines, where we can clearly identify the jump at $\phi=\pi$ when $\varphi=0$. We see that for $\varphi\neq 0$, the discontinuity in the supercurrent contribution corresponding to $k_{y}=0$ is shifted to $\phi\neq\pi$. Including all $k_{y}$-channels preserves the signature of the zero mode.
The case $\varphi=0$ (see plots in black in Fig. 4) is akin the fully gapped 2D topological superconductors with many channels, the hybridization of the propagating Majorana modes can be described by an effective Dirac Hamiltonian with a $\phi$-dependent mass. The Andreev levels associated to the hybridization of these states with a tunneling amplitude ${\cal T}_{\rm J}\propto t_{\rm J}$ have energies $\varepsilon_{k_{y}}(\phi)=\pm\sqrt{(vk_{y})^{2}+{\cal T}_{\rm J}^{2}\cos^{2}(\phi/2)}$ [Ruiz2022Nov]. For $k_{y}=0$, because of the existence of Majorana zero modes, these two states cross at $\phi=\pi$ and $J(\phi)$ has a jump at the value at that value of the phase bias. This is, in turn, identical to the behavior of 1D systems with the magnetic field perpendicular to the SOC [Lutchyn2010Aug, Oreg2010Oct, Kwon2004Feb, Pientka2013Feb]. The other edge-channels with $k_{y}\neq 0$ have a smooth non-sinusoidal behavior. In contrast, the $k_{y}$ channels corresponding to the bulk modes have a dispersion relation of the form $\varepsilon_{k_{y}}(\phi)=\pm E_{{\rm J},k_{y}}\cos(\phi)$, being $E_{{\rm J},k_{y}}$ a characteristic Josephson energy and ontribute to the CPR as $\propto\sin(\phi)$. The total $J(\phi)$, results from the contribution of all the Andreev states with negative energy. Therefore, the overall shape and features of the CPR are determined by the relative spectral weight of the edge modes among the transverse channels, and by the degree of their hybridization compared to the transparency for the bulk states.
For other orientations of $\varphi\neq 0$ within the topological phase (see plots in red and blue in 4), the level crossing of the component $k_{y}=0$, hence the jump in the CPR, occurs at different values of $\phi$ as in the case of 1D systems [Spanslatt2018Aug, Aligia2020Dec, Cayao2024Feb]. For orientations $\varphi$ beyond the condition of Eq. (6) for the boundary of the topological phase, the response is mainly dominated by a sinusoidal component for all the modes as in a non-topological superconductor (grey curve).
In all the cases, a response with $J(\phi=0)\neq 0$ is observed for $\varphi\neq 0$, typical of the anomalous Josephson effect. The amplitude of the critical current is also different for different signs of $\phi$, which is the characteristic feature of the Josephson diode effect [Reynoso2008Sep, BaumgartnerPRL2021, BaumgartnerNature2022] observed in the non-topological phase.
@@
Conclusions. We have shown the existence of a new topological phase in a 2D proximity-induced superconductor with Rashba spin-orbit coupling. This phase develops when a magnetic field is applied in-plane and is characterized by antichiral Majorana modes propagating along the edges perpendicular to the magnetic field. This phase takes place for a range of parameters (chemical potential and magnetic field) where the Bogoliubov-de Gennes spectrum has a single pair of Bogoliubov Fermi surfaces. We have shown that signatures of these modes can be identified in the behavior of the current-phase relation of a wide Josephson junction with the edges states oriented as sketched in Fig. 4. The topological phase persists for a range of angles close to this configuration. Other experimental signatures of these modes could be found in the heat transport [Read2000Apr, Nomura2012Jan, Sumiyoshi2013Jan], which is the counterpart to light propagation observed in photonic crystals with antichiral topological modes [Zhou2020Dec, Chen2022Jun, Hu2024Nov]. Although these experiments are challenging in superconductors, they have been implemented to study heat propagation along topological edge states in spin liquids [Kasahara2018Jul] and the quantum Hall state [Granger2009Feb, Altimiras2012Jul].
@@
The experimental realization of this phase is possible in various material platforms. Natural candidates are Al/InAs heterostructures to confine a 2D electron gas with spin orbit and proximity-induced superconductivity. In these systems, high-transparency Josephson junctions have been demonstrated, exhibiting anomalous Josephson effect and diode effect, while supporting large in-plane magnetic fields ($\sim$1.5T) [BaumgartnerPRL2021, BaumgartnerNature2022]. The typical electron density of the 2D gas is, albeit, not consistent with the range of chemical potentials for the topological phase. Nevertheless, a strategy can be developed to control it by using two quantum wells: one of them buried deep in the heterostructure to be used as a back-gate and the other one very shallow to be proximitized by the thin superconducting film. The gate tunability would allow to study the progressive effect of the emergence of pairs of Bogoliubov Fermi surfaces until the conditions for the topological phase are reached. Another potential platform is a two-dimensional magnetic topological insulator like that studied in Ref. [Mandal2024Jun].
@@
I Edge states in a disk
The spatial probability distribution of the lowest- positive-energy edge states in a disk is shown in Fig. 1. The finite size of the sample and the lattice discretization plays a role similar to disorder and generate interference patterns. However, the edge states can still be identified.
Figure S1: Disk symmetry: Spatial probability distribution of the lowest positive energy eigenstate $E\approx 0$ for an square lattice with $N=100000$ sites embeded in a circular shape with radius $R=90$ sites, $V=(1.1\Delta_{0},0,0)$, $E/t=1.7\times 10^{-5}$. The rest of the parameters are the same as in the main text.
II Details of the derivation of the effective low-energy Hamiltonian